# Using Dynamic Safety Margins as Control Barrier Functions

### v2

@@ 5 Lyapunov-based Dynamic Safety Margins
So, we must have $\Gamma(\boldsymbol{\mathbf{v}})=V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\Gamma^{*}(\boldsymbol{\mathbf{v}})$. If $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$, then $\boldsymbol{\mathbf{x}}=\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ because $\boldsymbol{\mathbf{v}}\in\mathcal{V}$, and we have a contradiction. Otherwise, if $\Gamma^{*}(\boldsymbol{\mathbf{v}})=V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})>0$, recall that $\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ is open in $\mathbb{X}$. Therefore, there exists an open set $\mathcal{N}\subset\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ with $\boldsymbol{\mathbf{x}}\in\mathcal{N}$. Furthermore, since $\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\setminus\{\bar{x}(\boldsymbol{\mathbf{v}})\}$ is a regular point of $V(\cdot,\boldsymbol{\mathbf{v}})$ by Assumption 4, we have that $\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}\neq 0$. So, there exists a sufficiently small scalar $\lambda>0$ such that $\tilde{\boldsymbol{\mathbf{x}}}\triangleq\boldsymbol{\mathbf{x}}-\lambda\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}^{\top}\in\mathcal{N}\subset\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. Furthermore, this point achieves lower value in $V$. That is, $V(\tilde{\boldsymbol{\mathbf{x}}},\boldsymbol{\mathbf{v}})<V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. Since $\boldsymbol{\mathbf{v}}\in\mathcal{V}$, we have that $\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. Thus, $\Gamma^{*}(\boldsymbol{\mathbf{v}})=\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. So, $\Gamma^{*}(\boldsymbol{\mathbf{v}})=V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ is a contradiction because
$\Gamma^{*}(\boldsymbol{\mathbf{v}})=\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq V(\tilde{\boldsymbol{\mathbf{x}}},\boldsymbol{\mathbf{v}})<V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),$
and we must have that $\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}$.
@@
To prove property (19c), let $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\times\mathbb{V}$ be such that $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$. By (31c),
$\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=-\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0.$ (39)
To prove the final property (19d), let $\boldsymbol{\mathbf{v}}\in\mathbb{V}$ and $\epsilon>0$ be given. If $\boldsymbol{\mathbf{v}}\notin\mathcal{V}$, then $\Gamma(\boldsymbol{\mathbf{v}})<0$ and $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$ for all $\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. Thus, $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})<0$ for all $\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ and we have that $\mathcal{C}_{\boldsymbol{\mathbf{v}}}^{\epsilon}=\emptyset$ which is trivially closed in $\mathbb{X}$. Otherwise, if $\boldsymbol{\mathbf{v}}\in\mathcal{V}$, let $\{\boldsymbol{\mathbf{x}}_{n}\}\subset\mathcal{C}_{\boldsymbol{\mathbf{v}}}^{\epsilon}$ be a sequence that converges to $\tilde{\boldsymbol{\mathbf{x}}}\in\mathbb{X}$. Define the map $\zeta:\boldsymbol{\mathbf{x}}\mapsto\Gamma(\boldsymbol{\mathbf{v}})-V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. By continuity of $\zeta$ on $\mathbb{X}$, we have that the set $\mathcal{S}=\{\boldsymbol{\mathbf{x}}\in\mathbb{X}:\zeta(\boldsymbol{\mathbf{x}})\geq\epsilon\}$ is closed in $\mathbb{X}$. Since $\{\boldsymbol{\mathbf{x}}_{n}\}\subset\mathcal{C}_{\boldsymbol{\mathbf{v}}}^{\epsilon}\subset\mathcal{S}$, we have that $\tilde{\boldsymbol{\mathbf{x}}}\in\mathcal{S}$. It remains to show that $\tilde{\boldsymbol{\mathbf{x}}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. Begin by noting that $\zeta(\tilde{\boldsymbol{\mathbf{x}}})=\Gamma(\boldsymbol{\mathbf{v}})-V(\tilde{\boldsymbol{\mathbf{x}}},\boldsymbol{\mathbf{v}})\geq\epsilon$ because $\tilde{\boldsymbol{\mathbf{x}}}\in\mathcal{S}$. Thus,
$\min\big{(}\Gamma^{*}(\boldsymbol{\mathbf{v}}),\overline{\Gamma}(\boldsymbol{\mathbf{v}})\big{)}\geq\Gamma(\boldsymbol{\mathbf{v}})\geq V(\tilde{\boldsymbol{\mathbf{x}}},\boldsymbol{\mathbf{v}})+\epsilon.$
For a contradiction, assume $\tilde{\boldsymbol{\mathbf{x}}}\in\mathcal{D}^{c}_{\boldsymbol{\mathbf{v}}}=\mathbb{X}\setminus\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. If $\tilde{\boldsymbol{\mathbf{x}}}\in\partial\mathcal{D}_{\boldsymbol{\mathbf{v}}}\subset\mathcal{D}^{c}_{\boldsymbol{\mathbf{v}}}$, then we have that
$\overline{\Gamma}(\boldsymbol{\mathbf{v}})=\inf_{\boldsymbol{\mathbf{x}}\in\partial\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq V(\tilde{\boldsymbol{\mathbf{x}}},\boldsymbol{\mathbf{v}})+\epsilon>V(\tilde{\boldsymbol{\mathbf{x}}},\boldsymbol{\mathbf{v}}),$
which is a contradiction. On the other hand, if $\tilde{\boldsymbol{\mathbf{x}}}\in\operatorname{Int}\left(\mathcal{D}^{c}_{\boldsymbol{\mathbf{v}}}\right)\subset\mathcal{D}^{c}_{\boldsymbol{\mathbf{v}}}$, there exists an open set $\mathcal{N}\subset\operatorname{Int}\left(\mathcal{D}^{c}_{\boldsymbol{\mathbf{v}}}\right)$ with $\tilde{\boldsymbol{\mathbf{x}}}\in\mathcal{N}$. By convergence of $\{\boldsymbol{\mathbf{x}}_{n}\}$ to $\tilde{\boldsymbol{\mathbf{x}}}$, there exists $N\in\mathbb{N}$ such that $\boldsymbol{\mathbf{x}}_{N}\in\mathcal{N}\subset\operatorname{Int}\left(\mathcal{D}^{c}_{\boldsymbol{\mathbf{v}}}\right)\subset\mathcal{D}^{c}_{\boldsymbol{\mathbf{v}}}$. This is a contradiction because $\boldsymbol{\mathbf{x}}_{N}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. So, we must have that $\tilde{\boldsymbol{\mathbf{x}}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{S}=\mathcal{C}_{\boldsymbol{\mathbf{v}}}^{\epsilon}$. It follows that $\mathcal{C}_{\boldsymbol{\mathbf{v}}}^{\epsilon}$ is closed in $\mathbb{X}$.
@@
If $\min\big{(}\Gamma^{*}(\boldsymbol{\mathbf{v}}),\overline{\Gamma}(\boldsymbol{\mathbf{v}})\big{)}$ is differentiable, taking $\Gamma(\boldsymbol{\mathbf{v}})=(1-\lambda)\min\big{(}\Gamma^{*}(\boldsymbol{\mathbf{v}}),\overline{\Gamma}(\boldsymbol{\mathbf{v}})\big{)}$, with $\lambda\in(0,1)$ small can make $\mathcal{C}$ compact.
@@
When multiple constraints are considered, the stability threshold value $\overline{\Gamma}(\boldsymbol{\mathbf{v}})$ is shared among them. Due to this, only one of the $\Gamma_{i}$ needs to satisfy $\Gamma_{i}(\boldsymbol{\mathbf{v}})\leq\overline{\Gamma}(\boldsymbol{\mathbf{v}})$ as in Theorem 5.26. However, every $\Gamma_{i}$ must respect $\Gamma_{i}(\boldsymbol{\mathbf{v}})\leq\Gamma_{i}^{*}(\boldsymbol{\mathbf{v}})$.
Remark 5.30.
When choosing the initial virtual reference $\boldsymbol{\mathbf{v}}$ in the augmented system (22), it is worth noting that any $\boldsymbol{\mathbf{v}}_{0}\in\mathcal{V}$ such that $(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}}_{0})\in\mathcal{C}$ is valid. A desirable choice is

### v3

@@ 4 Lyapunov-based Dynamic Safety Margins
which implies that $\Gamma^{*}(\boldsymbol{\mathbf{v}})\leq\Gamma(\boldsymbol{\mathbf{v}})\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. So, we must have $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\Gamma^{*}(\boldsymbol{\mathbf{v}})$. If $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$, then $\boldsymbol{\mathbf{x}}=\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ because $\boldsymbol{\mathbf{v}}\in\mathcal{V}$, and we have a contradiction. Otherwise, if $\Gamma^{*}(\boldsymbol{\mathbf{v}})=V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})>0$, recall that $\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ is open in $\mathbb{R}^{n}$. Therefore, there exists an open set $\mathcal{N}\subset\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ with $\boldsymbol{\mathbf{x}}\in\mathcal{N}$. Furthermore, since $\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\setminus\{\bar{x}(\boldsymbol{\mathbf{v}})\}$ is a regular point of $V(\cdot,\boldsymbol{\mathbf{v}})$ by Assumption 4, we have that $\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}\neq 0$. So, there exists a sufficiently small scalar $\delta>0$ such that $\hat{\boldsymbol{\mathbf{x}}}=\boldsymbol{\mathbf{x}}-\delta\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}^{\top}\in\mathcal{N}\subset\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. Furthermore, this point achieves lower value in $V$. That is, $V(\hat{\boldsymbol{\mathbf{x}}},\boldsymbol{\mathbf{v}})<V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. So, $\Gamma^{*}(\boldsymbol{\mathbf{v}})=V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ is a contradiction because
$\Gamma^{*}(\boldsymbol{\mathbf{v}})=\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq V(\hat{\boldsymbol{\mathbf{x}}},\boldsymbol{\mathbf{v}})<V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}).$
To prove property (22b), note that since $\tilde{\mathcal{C}}\subset\mathcal{X}\times\mathcal{V}$, and $\mathcal{X}\times\mathcal{V}$ is compact in $\mathbb{R}^{n}\times\mathbb{R}^{l}$, we must only show that $\tilde{\mathcal{C}}$ is closed in $\mathbb{R}^{n}\times\mathbb{R}^{l}$. Let $\{(\boldsymbol{\mathbf{x}}_{n},\boldsymbol{\mathbf{v}}_{n})\}\subset\tilde{\mathcal{C}}$ be a sequence that converges to $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathbb{R}^{n}\times\mathbb{R}^{l}$. Since $\mathcal{X}\subset\mathbb{R}^{n}$ and $\mathcal{V}\subset\mathbb{R}^{l}$ are closed, it follows that $\boldsymbol{\mathbf{x}}\in\mathcal{X}$ and $\boldsymbol{\mathbf{v}}\in\mathcal{V}$. Now we argue $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$. For a contradiction, assume there is $i\in\{1,2\}$ such that $\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})<0$. By continuity of $\Delta_{i}$ on $\mathbb{R}^{n}\times\mathbb{R}^{l}$, there exists an open neighborhood $\tilde{\mathcal{N}}\subset\mathbb{R}^{n}\times\mathbb{R}^{l}$ of $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ wherein $\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})<0$. By convergence of $\{(\boldsymbol{\mathbf{x}}_{n},\boldsymbol{\mathbf{v}}_{n})\}$ to $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$, there exists $N\in\mathbb{N}$ such that $(\boldsymbol{\mathbf{x}}_{N},\boldsymbol{\mathbf{v}}_{N})\in\tilde{\mathcal{N}}$, which raises a contradiction because $(\boldsymbol{\mathbf{x}}_{N},\boldsymbol{\mathbf{v}}_{N})\in\tilde{\mathcal{C}}$ implies that $\Delta(\boldsymbol{\mathbf{x}}_{N},\boldsymbol{\mathbf{v}}_{N})\geq 0$. It remains to show that $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{D}}$. Begin by noting that for all $n\in\mathbb{N}$, $(\boldsymbol{\mathbf{x}}_{n},\boldsymbol{\mathbf{v}}_{n})\in\tilde{\mathcal{D}}$ implies that $V(\boldsymbol{\mathbf{x}}_{n},\boldsymbol{\mathbf{v}}_{n})\geq\gamma_{1}\big{(}\|\boldsymbol{\mathbf{x}}_{n}-\bar{x}(\boldsymbol{\mathbf{v}}_{n})\|\big{)}$. It follows by continuity of all functions involved in the inequality and by the order limit theorem that $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq\gamma_{1}\big{(}\|\boldsymbol{\mathbf{x}}-\bar{x}(\boldsymbol{\mathbf{v}})\|\big{)}$. For a contradiction, let us assume that $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\partial\tilde{\mathcal{D}}$, which implies that $\boldsymbol{\mathbf{x}}\in\partial\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ and, therefore, $\overline{\Gamma}(\boldsymbol{\mathbf{v}})\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. However, recall that we have shown $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$. Thus, $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq(1-\epsilon)\overline{\Gamma}(\boldsymbol{\mathbf{v}})$. This yields
$\frac{1}{1-\epsilon}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq\bar{\Gamma}(\boldsymbol{\mathbf{v}})\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),$
which is a contradiction because $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq\gamma_{1}\big{(}\|\boldsymbol{\mathbf{x}}-\bar{x}(\boldsymbol{\mathbf{v}})\|\big{)}>0$. So, we must have that $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{D}}$ and we conclude that $\tilde{\mathcal{C}}$ is closed in $\mathbb{R}^{n}\times\mathbb{R}^{l}$, and therefore compact.
@@
To prove property (22c), let $i\in\{1,2\}$ be given and pick $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{D}}$ such that $\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$. Then, by (33b),
$\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=-\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0.$ (38)
The following corollary highlights that for Lyapunov-based DSMs, any $\alpha\in\mathscr{K}_{\infty}$ can be used for the DSM-CBF policy $\beta$ given in (31). This result constitutes a step away from existence-type proofs common in the CBF literature and towards implementation.
Corollary 4.27.
If $\Delta$ is a Lyapunov-based DSM as in (37), then $\forall\alpha\in\mathscr{K}_{\infty}$, $\forall(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, $\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\neq\emptyset$.
Proof 4.28.
Let $\alpha\in\mathscr{K}_{\infty}$ and $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$ be given. Recall that $\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{U}$ and note that for any $i\in\{1,2\}$,
$\dot{\Delta}_{i}\big{(}\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}},\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}=-\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0\geq-\alpha\big{(}\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}.$
Therefore, $\big{(}\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}\in\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$.
@@
When multiple constraints are considered, it is usually simpler to define multiple safety threshold values $\Gamma^{*}_{i}$ and stack them into the form of (37). As long as each one is differentiable, Theorem 4.25 can still be applied to obtain a DSM $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ with $n_{c}>2$. In the event either $\Gamma^{*}_{i}$ or $\overline{\Gamma}$ are not continuously differentiable, any continuously differentiable, lower-bounding function can be used in its place.
@@
When choosing the initial virtual reference $\boldsymbol{\mathbf{v}}$ in the augmented system (24), it is worth noting that any $\boldsymbol{\mathbf{v}}_{0}\in\mathcal{V}$ such that $(\boldsymbol{\mathbf{x}}_{0},\boldsymbol{\mathbf{v}}_{0})\in\tilde{\mathcal{C}}$ is valid. A desirable choice is
@@
Remark 4.31.
The proposed method effectively transforms the problem of finding a CBF into that of finding a Lyapunov function. The advantage in doing so is that the latter problem has a much richer history, which can be leveraged to synthesize Lyapunov-based DSMs using both analytical [30, 31, 32] and numerical [34] tools.
5 DSM-CBF Synthesis Approach Summary
In this manuscript, we have shown that dynamic safety margins (DSMs), key components in the explicit reference governor ERG, are (vector-valued) CBFs for an augmented system obtained by concatenating the state and reference vectors. This insight allows the direct use of approaches that compute DSMs to synthesize CBFs. Existing ERG works like [23] outline different approaches to construct DSMs. This manuscript studied the Lyapunov-based approach, which requires finding a reference-dependent Lyapunov function (33) for the prestabilized system and evaluating the safety (34) and stability (36) threshold values. These ingredients can then be combined to form a DSM. Once a DSM is obtained, a safe constrained control policy that minimally adjusts the nominal controller can be readily implemented by solving (31). The following examples illustrate using the Lyapunov-based approach to construct DSM-CBFs and implementing (31). Additional examples can be found in [39], where the approach presented in this paper is specialized to Euler–Lagrange systems.